OpenStudy (anonymous):

Write the trigonometric expression as an algebraic expression containing u and v. sin(tan^-1 (u) - tan^-1 (v))

OpenStudy (anonymous):

Using sin(A-B)= sinAcosB - cos AsinB Sin(tan^-1 u- tan-1 v)= sin(tan^-1u)cos(tan^-1v)-cos(tan^-1u)s… Now if an angle has a tanx= u/1, then the hypotenuse is sqr(u^2+1), so sin x = u/sqr(u^2+1) and cosx= 1/sqr(u^2+1) sin(tan^-1u)cos(tan^-1v)-cos(tan^-1u)s… [u/sqr(u^2+1)][1/sqr(v^2+1)]- [1/sqr(u^2+1)][v/sqr(v^2+1)]...... = (u-v)/[sqr(u2+1)][sqr(v^2+1)] Now you can simplify.

OpenStudy (primeralph):

Stuff like this makes me dizzy.

OpenStudy (primeralph):

@natalia0207 Do you follow?

OpenStudy (primeralph):